OpenStudy (anonymous):

Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3,-4). Help?

OpenStudy (anonymous):

I don't want the answer. I would like someone to actually show me how to work through this! I'm not good with this section in Geometry

OpenStudy (tkhunny):

Here is the point-slope form. \[(y-y_{0})=m(x-x_{0})\] Substitute and simplify!

OpenStudy (anonymous):

But they give me only one pair of points so how exactly do I go about doin this?

OpenStudy (tkhunny):

You have three things 1) A Point 2) A slope 3) The point-slope form That is all you need. Substitute the points coordiantes into y0 and x0 and the slope in m. You'll be done.

OpenStudy (anonymous):

So, would it be (0,-4)=-3(3,0) I'm still a little confused.

hero (hero):

If you're going to write the point-slope form, at least write it this way: \[y - y_1 = m(x - x_1)\] You are given \(m = -3\) and \((x_1, y_1) = (3,-4)\). Only substitute what you are given into the equation.

OpenStudy (anonymous):

(0,-4)=-3 (3,0) right?

hero (hero):

Nowhere in the problem are you given (x,y) = (0,0). If you are not given something, don't substitute it. Not every single variable needs to be substituted.

hero (hero):

And also, the formula does not tell you to include any commas.

OpenStudy (anonymous):

oh! my bad! But I do not understand what I would put for x and y? would it be y-(-4)=-3(x-3)?

hero (hero):

You have it

OpenStudy (anonymous):

Since in the left side it is two negatives, would I turn it into a positive? or leave it?

hero (hero):

y-(-4) = -3(x-3) is correct, however it should be simplified to y + 4 = -3(x-3)

hero (hero):

Good job

OpenStudy (anonymous):

Thank you!! Now I understand how to do the rest!(: